# thanks: Equal contributionthanks: Equal contribution

### v1

@@ Abstract
Interpreting multi-messenger signals from neutron stars and black holes requires reliable general-relativistic magnetohydrodynamics (GRMHD) simulations across rapidly evolving high-performance-computing platforms, yet key algorithms are routinely rewritten within infrastructure-specific numerical-relativity codes, hindering verification and reuse. We present the General Relativistic Hydrodynamics Library (GRHayL), a modular, infrastructure-agnostic GR(M)HD library providing conservative-to-primitive recovery, reconstruction, flux/source and induction operators, equations of state, and neutrino leakage through an intuitive interface. GRHayL refactors and extends the mature IllinoisGRMHD code into reusable pointwise and stencil-wise kernels, enabling rapid development and cross-code validation in diverse frameworks, while easing adoption of new microphysics and future accelerators. We implement the same kernels in the Einstein Toolkit (Carpet and CarpetX) and BlackHoles@Home, demonstrating portability with minimal duplication. Validation combines continuous-integration unit tests with cross-infrastructure comparisons of analytic GRMHD Riemann problems, dynamical Tolman–Oppenheimer–Volkoff evolutions, and binary neutron-star mergers, showing comparable or improved behavior over legacy IllinoisGRMHD and established Einstein Toolkit codes.
@@ I Introduction
Over the past decade, gravitational wave (GW) astronomy has experienced rapid growth in scientific output, starting with the first direct detection of GWs [GW150914] and now encompassing over 200 compact binary coalescence detections [GW151226, GW170104, GW170608, GW170814, GW170817, GW190412, GW190521, GW190521_analysis, GW190814, GW200105, LIGO_catalog1, LIGO_catalog2, LIGO_catalog3, LIGO_catalog4, LIGO_catalog5]. Among these first observations, perhaps the most impactful was GW170817/GRB170817A/AT2017gfo [GW170817], the only multi-messenger binary neutron star (BNS) merger detection to date. The GW signal was accompanied by observations of gamma rays and months-long emission throughout much of the rest of the electromagnetic (EM) spectrum [GW170817_GRB, Goldstein_2017, Tanvir_2017]. This event is expected to be just the first in GW-anchored multi-messenger astronomy, which combines data from GW signals with data from coincident EM and, for sufficiently nearby sources, possibly neutrino detections to investigate some of the most extreme events in the cosmos. Looking forward, iterative improvements to the Advanced LIGO detectors, as well as future ground-based [CosmicExplorer1, CosmicExplorer2, punturo2010einstein] and space-based [LISA_2023, Luo_2016] GW detectors, will collectively provide remarkably improved detection capabilities, triggering neutrino and electromagnetic follow-ups across the spectrum.
@@
Beyond BNS mergers, multi-messenger observations of black hole-neutron star (BHNS) systems, circumbinary disks, core-collapse supernovae, and other compact-object environments are poised to greatly expand our understanding of high-energy astrophysical systems. However, these insights are gained by comparing observations with theoretical models, which in turn will be limited by how faithfully they track the underlying physics.
@@
Generating self-consistent models of important multi-messenger sources is exceedingly challenging, as it in principle requires fully solving the equations of general relativity (GR) for the spacetime, combined with a magnetohydrodynamic (MHD) treatment of turbulent magnetized plasmas, nuclear equations of state (EoSs), neutrino and EM radiation transport, r-process nucleosynthesis via nuclear reaction networks, and detailed EM ray tracing. Fully self-consistent modeling of all of these processes in a single three-dimensional simulation exceeds the computational resources currently available to the community, not to mention that the relevant length and time scales span several orders of magnitude, making even one such calculation prohibitively expensive. As NR codes are extended to incorporate additional microphysics and radiation effects, the underlying frameworks have historically counterbalanced the growing computational cost by taking advantage of advances in central processing unit (CPU) architectures. However, just as the revolution in multi-messenger astronomy has greatly increased the importance of improving and extending the physical realism of NR models, the rapid shift towards heterogeneous high-performance computing (HPC) systems and reliance on accelerators like graphics processing units (GPUs)—heavily driven by the advent of artificial intelligence and machine learning—now require that code frameworks be upgraded to take full advantage of state-of-the-art HPC resources.
@@
Transforming existing CPU-tuned NR frameworks to be performant and scalable on heterogeneous CPU/GPU HPC clusters typically requires a substantial redesign of the underlying software infrastructure. As a result, significant effort has been expended in recent years to migrate existing simulation software to computational frameworks that explicitly support heterogeneous architectures and offer performance portability across devices. Many groups are working towards extending their in-house infrastructures [Ishi_2019] and/or developing entirely new ones [ET_2023_11] that build on performance-portable programming models or modern AMR infrastructures in order to fully exploit current and future HPC resources.
@@
Two notable examples of technologies enabling this transition are Kokkos [kokkos] and AMReX [amrex], which play complementary roles. Kokkos is a general-purpose performance-portability library that provides an abstraction layer for expressing parallelism and memory access patterns in a way that can be mapped efficiently onto diverse heterogeneous HPC architectures. In the NR community, Kokkos has been used to extend Athena++ [Stone_2020, GR_Athena] to be GPU-capable. The Parthenon framework [Grete:2023] further develops the original Athena++ code into a more general framework for multiphysics code development on top of Kokkos. Several other projects connected to the Los Alamos National Laboratory also provide open-source components built on Kokkos that are directly relevant to NR, such as the portability abstraction library Ports-of-Call [PortsOfCall], the tabulated data interpolation library Spiner [miller2022notquite, Miller2022], and the EoS library Singularity EOS [miller2024singularity].
@@
In parallel, AMReX provides a general framework for implementing block-structured adaptive mesh refinement (AMR) codes on heterogeneous HPC systems. It supplies the mesh-management, load-balancing, and parallelization infrastructure needed to build scalable AMR applications on modern CPU/GPU clusters, and has been used by several groups to develop NR and NR-adjacent codes for these systems. The Einstein Toolkit [ET_2023_11], primarily built on the Cactus Framework [Cactus] and the Carpet AMR driver, is one example of an existing code base undergoing such a modernization. Recently, the new AMReX-based CarpetX code was developed and released in the Einstein Toolkit as a GPU-capable replacement for Carpet. Another notable example is GRChombo [Clough_2015, Andrade2021], whose developers are in the process of creating GRTeclyn and porting the science code to the new infrastructure.
@@
As all these efforts show, new infrastructures are required to fully leverage current and future developments in HPC and numerical techniques. Thus, in order to keep pace with these changes, there is a clear and urgent need to minimize the time and effort required to migrate NR codes between infrastructures. The current paradigm often requires significant code rewrites or refactoring, which is undesirable, as most GRMHD algorithms are independent of the infrastructure.
@@
Several groups have released code packages that partially address this issue. For instance, several open-source codes provide functions for conservative-to-primitive recovery schemes [Noble:2005gf, Siegel:2017jug] and interpolators for finite-temperature EoS tables [OConnor:2009iuz, Schneider:2017tfi]. While these codes require some work to incorporate into a new code, they do provide core algorithms in a reasonably self-contained form. Additionally, several libraries have been developed to provide implementations of EoS [miller2022notquite, Miller2022, miller2024singularity], conservative-to-primitive recovery schemes [Kastaun_2021, Roland_2024], and neutrino physics [OConnor:2014sgn, Cheong:2023fgh].
@@
These groups have demonstrated both the success of and demand for infrastructure- and hardware-agnostic implementations of the core GRMHD algorithms needed for NR simulations. To more comprehensively address the needs of GRMHD simulations, we introduce the General Relativistic Hydrodynamics Library (GRHayL). This library contains all the core elements required for a GRMHD simulation and represents a refactoring and extension of the IllinoisGRMHD code [Duez_IGM, Etienne_IGM]—including the addition of tabulated EoS and neutrino leakage support in the forked code by Werneck et al. [Werneck_IGM]—into a modular, infrastructure-agnostic library. GRHayL provides nearly all of the code required to develop an IllinoisGRMHD-like code via pointwise or stencil-wise library functions, allowing for quicker code development, easier cross-code and cross-infrastructure collaboration, and cleaner implementations of GRMHD codes. There is also ongoing work to add full GPU support.
@@
This paper is organized as follows. Section II discusses the individual code modules within the library and provides an overview of the available functionality. Section III lists all the functional changes and improvements we have made to the IllinoisGRMHD code while developing the library. Section IV describes the codes we have developed using GRHayL as well as the infrastructures themselves, in preparation for Section V, which compares data from the original IllinoisGRMHD code and various new implementations using GRHayL. Finally, in Section VI we summarize the library features and discuss planned extensions of GRHayL’s capabilities and implementations in additional infrastructures. In addition, we include two appendices. Appendix A briefly describes the GRMHD evolution equations and our choice of variables used for evolving them. Appendix B details the automated, function-level continuous integration testing that we perform to validate the library for each change made to the code repository.
@@ II Library Features
The core part of the library—the “chalice” of GRHayL—is the glue that ties all the separate modules together. This must be as minimal as possible to meet the aforementioned requirements of the library. GRHayL’s core contains definitions of C structs and simple helper functions for packing and unpacking data to and from these structs, as well as performing basic operations dependent purely on struct quantities (e.g., computing $g_{\mu\nu}$ and $g^{\mu\nu}$ from the ADM variables).
@@ II.1 EOS Gem
The EOS gem provides EoS functionality within GRHayL. It currently supports hybrid piecewise polytropic EoSs and includes infrastructure for fully tabulated EoSs. Tabulated EoS support is presently under development and is not yet intended for production simulations. Many other gems depend on functions provided by the EOS gem, but this dependency has been abstracted to allow users to implement their own EoS driver without changing the other gems. The EoS information is conveyed primarily via an EOS struct, in addition to several function pointers that are set during initialization. Users are free to swap the EOS functions with their own by setting these function pointers. Further, other EoS types or implementations can be easily added to the gem. For example, providing support for the Singularity EOS library would simply require defining the proper wrapper functions that convert the expected arguments for the GRHayL functions and making the appropriate call to Singularity EOS.
A hybrid EoS [1993A&A...268..360J] assumes that the pressure and specific internal energy can be split into cold and thermal components, i.e.,
@@
In GRHayL, the cold pressure is represented by a piecewise polytrope (see, e.g., [Read:2008iy]),
@@
where $\Gamma_{\mathrm{th}}$ is a constant parameter that determines the conversion efficiency of kinetic to thermal energy at shocks. The EOS gem additionally supports a special case of the hybrid EoS called the simple EoS. [^2] This is often referred to as the “ideal fluid”, “ideal gas”, or “Gamma-law” EoS. In this case, the EoS contains a single piece with $\Gamma_{0}=\Gamma_{\mathrm{th}}$, resulting in $P=(\Gamma_{0}-1)\rho\epsilon$.
@@ II.2 Con2Prim Gem
The Con2Prim gem provides several methods for recovering the primitive variables $\bm{P}$ from the conservative variables $\bm{C}$ (defined in Eqs. 12 and 16, respectively). Because of the nonlinearity of the equations, a root-finding method such as Newton–Raphson is required. In addition, this gem provides several helper functions to compute $\bm{C}(\bm{P})$, to produce initial guesses for the primitives, and to enforce bounds on $\bm{C}$ [Faber1, Faber2] and $\bm{P}$.
The gem includes several primitive recovery routines, including some by Noble et al. [Noble:2005gf, Noble:2008tm], Font et al. [Font_2000, Etienne_2012_3], Palenzuela et al. [Palenzuela_2015], and ongoing efforts to support Newman & Hamlin [Newman_2014] (see also [Werneck_IGM] for an entropy variation of the last two). Routines are named based on the first author of their originating paper and the dimensionality of the method (e.g., Noble2D, Palenzuela1D, etc.). The gem also provides several useful diagnostics, including the number of iterations a routine took to converge to the solution and whether or not backup routines were used.
The routines are adapted from origIGM and from [Siegel:2017jug], noting that origIGM originally adapted the Noble et al. routines from HARM3D [Noble:2005gf, Noble:2008tm] and HARM3D+NUC [Murguia-Berthier:2021tnt]. We summarize all available primitive recovery routines currently implemented in GRHayL in Table 1.
@@
Noble2D [Noble:2005gf]
Noble1D [Noble:2005gf]
Noble1D_entropy [Noble:2008tm]
Font1D [Font_2000, Etienne_2012_3]
Palenzuela1D [Palenzuela_2015]
Palenzuela1D_entropy [Werneck_IGM]
@@ II.3 Flux_Source Gem
The Flux_Source gem provides functions to compute the flux and source terms for the right-hand sides (RHSs) of the evolved GRMHD variables as described in Appendix A. These functions calculate the flux contributions to the RHSs using a finite-volume method, and the needed reconstructions can be performed using the Reconstruction gem. Flux_Source uses NRPy [Ruchlin:2018] to automatically generate optimized C code from the basic flux and source equations. The Flux_Source gem currently only provides the Harten, Lax, & van Leer (HLL) [Harten_1983] approximate Riemann solver, which has the form
@@
Here $c^{-}$ is the minimum characteristic speed, $c^{+}$ is the maximum characteristic speed, $F_{r,l}$ are the hydrodynamic fluxes defined in Eq. (20), and $U_{r,l}$ are the reconstructed values for the conserved variable $U$. The $r,l$ subscripts indicate fluid and spacetime quantities reconstructed at the right and left faces along some $i$-th flux direction. To compute the characteristic speeds, we follow the prescription given by [HARM], which approximates the general GRMHD dispersion relation as a quadratic in the wave speed. The wave speeds are then given by evaluating the quadratic equations

### v2

@@ Abstract
Interpreting multimessenger signals from neutron stars and black holes requires reliable general relativistic magnetohydrodynamics (GRMHD) simulations across rapidly evolving high-performance computing platforms, yet key algorithms are routinely rewritten within infrastructure-specific numerical relativity codes, hindering verification and reuse. We present the General Relativistic Hydrodynamics Library (GRHayL), a modular, infrastructure-agnostic GR(M)HD library providing conservative-to-primitive recovery, reconstruction, flux/source and induction operators, equations of state, and neutrino leakage through an intuitive interface. GRHayL refactors and extends the mature IllinoisGRMHD code into reusable point- and stencilwise kernels, enabling rapid development and cross-code validation in diverse frameworks, while easing adoption of new microphysics and future accelerators. We implement the same kernels in the Einstein Toolkit (Carpet and CarpetX) and BlackHoles@Home, demonstrating portability with minimal duplication. Validation combines continuous-integration unit tests with cross-infrastructure comparisons of analytic GRMHD Riemann problems, dynamical Tolman–Oppenheimer–Volkoff evolutions, and binary neutron star mergers, showing comparable or improved behavior over legacy IllinoisGRMHD and established Einstein Toolkit codes.
@@ I Introduction
Over the past decade, gravitational wave (GW) astronomy has experienced rapid growth in scientific output, starting with the first direct detection of GWs [3] and now encompassing over 200 compact binary coalescence detections [2, 6, 7, 8, 9, 11, 12, 14, 13, 16, 10, 15, 18, 17, 1]. Among these first observations, perhaps the most impactful was GW170817/GRB170817A/AT2017gfo [9], the only multimessenger binary neutron star (BNS) merger detection to date. The GW signal was accompanied by observations of $\gamma$ rays and months-long emission throughout much of the rest of the electromagnetic (EM) spectrum [5, 57, 110]. This event is expected to be just the first in GW-anchored multimessenger astronomy, which combines data from GW signals with data from coincident EM and, for sufficiently nearby sources, possibly neutrino detections to investigate some of the most extreme events in the cosmos. Looking forward, iterative improvements to the Advanced LIGO detectors, as well as future ground-based [39, 4, 93] and space-based [19, 71] GW detectors, will collectively provide remarkably improved detection capabilities, triggering neutrino and electromagnetic follow-ups across the spectrum.
@@
Beyond BNS mergers, multimessenger observations of black hole–neutron star systems, circumbinary disks, core-collapse supernovae, and other compact-object environments are poised to greatly expand our understanding of high-energy astrophysical systems. However, these insights are gained by comparing observations with theoretical models, which in turn will be limited by how faithfully they track the underlying physics.
@@
Generating self-consistent models of important multimessenger sources is exceedingly challenging, as it, in principle, requires fully solving the equations of general relativity for the spacetime, combined with a magnetohydrodynamic (MHD) treatment of turbulent magnetized plasmas, nuclear equations of state (EoS), neutrino and EM radiation transport, r-process nucleosynthesis via nuclear reaction networks, and detailed EM ray tracing. Fully self-consistent modeling of all of these processes in a single three-dimensional simulation exceeds the computational resources currently available to the community, not to mention that the relevant length and timescales span several orders of magnitude, making even one such calculation prohibitively expensive. As NR codes are extended to incorporate additional microphysics and radiation effects, the underlying frameworks have historically counterbalanced the growing computational cost by taking advantage of advances in central processing unit (CPU) architectures. However, just as the revolution in multimessenger astronomy has greatly increased the importance of improving and extending the physical realism of NR models, the rapid shift toward heterogeneous high-performance computing (HPC) systems and reliance on accelerators like graphics processing units (GPUs)—heavily driven by the advent of artificial intelligence and machine learning—now require that code frameworks be upgraded to take full advantage of state-of-the-art HPC resources.
@@
Transforming existing CPU-tuned NR frameworks to be performant and scalable on heterogeneous CPU/GPU HPC clusters typically requires a substantial redesign of the underlying software infrastructure. As a result, significant effort has been expended in recent years to migrate existing simulation software to computational frameworks that explicitly support heterogeneous architectures and offer performance portability across devices. Many groups are working toward extending their in-house infrastructures [62] and/or developing entirely new ones [102] that build on performance-portable programming models or modern adaptive mesh refinement (AMR) infrastructures in order to fully exploit current and future HPC resources.
@@
Two notable examples of technologies enabling this transition are Kokkos [112] and AMReX [116], which play complementary roles. Kokkos is a general-purpose performance-portability library that provides an abstraction layer for expressing parallelism and memory access patterns in a way that can be mapped efficiently onto diverse heterogeneous HPC architectures. In the NR community, Kokkos has been used to extend Athena++ [109, 35] to be GPU capable. The Parthenon framework [59] further develops the original Athena++ code into a more general framework for multiphysics code development on top of Kokkos. Several other projects connected to the Los Alamos National Laboratory also provide open-source components built on Kokkos that are directly relevant to NR, such as the portability abstraction library Ports-of-Call [78], the tabulated data interpolation library Spiner [76, 77], and the EoS library Singularity EOS [75].
@@
In parallel, AMReX provides a general framework for implementing block-structured adaptive mesh refinement (AMR) codes on heterogeneous HPC systems. It supplies the mesh-management, load-balancing, and parallelization infrastructure needed to build scalable AMR applications on modern CPU/GPU clusters, and has been used by several groups to develop NR and NR-adjacent codes for these systems. The Einstein Toolkit [102], primarily built on the Cactus framework [58] and the Carpet AMR driver, is one example of an existing code base undergoing such a modernization. Recently, the new AMReX-based CarpetX code was developed and released in the Einstein Toolkit as a GPU-capable replacement for Carpet. Another notable example is GRChombo [31, 20], whose developers are in the process of creating GRTeclyn and porting the science code to the new infrastructure.
@@
As all these efforts show, new infrastructures are required to fully leverage current and future developments in HPC and numerical techniques. Thus, in order to keep pace with these changes, there is a clear and urgent need to minimize the time and effort required to migrate NR codes between infrastructures. The current paradigm often requires significant code rewrites or refactoring, which is undesirable, as most general relativistic magnetohydrodynamics (GRMHD) algorithms are independent of the infrastructure.
@@
Several groups have released code packages that partially address this issue. For instance, several open-source codes provide functions for conservative-to-primitive recovery schemes [86, 108] and interpolators for finite-temperature EoS tables [88, 103]. While these codes require some work to incorporate into a new code, they do provide core algorithms in a reasonably self-contained form. Additionally, several libraries have been developed to provide implementations of EoS [76, 77, 75], conservative-to-primitive recovery schemes [68, 60], and neutrino physics [89, 29].
@@
These groups have demonstrated both the success of and demand for infrastructure- and hardware-agnostic implementations of the core GRMHD algorithms needed for NR simulations. To more comprehensively address the needs of GRMHD simulations, we introduce the General Relativistic Hydrodynamics Library (GRHayL). This library contains all the core elements required for a GRMHD simulation and represents a refactoring and extension of the IllinoisGRMHD code [38, 41]—including the addition of tabulated EoS and neutrino leakage support in the forked code by Werneck et al. [115]—into a modular, infrastructure-agnostic library. GRHayL provides nearly all of the code required to develop an IllinoisGRMHD-like code via point- or stencilwise library functions, allowing for quicker code development, easier cross-code and cross-infrastructure collaboration, and cleaner implementations of GRMHD codes. There is also ongoing work to add full GPU support.
@@
This paper is organized as follows. Section II discusses the individual code modules within the library and provides an overview of the available functionality. Section III lists all the functional changes and improvements we have made to the IllinoisGRMHD code while developing the library. Section IV describes the codes we have developed using GRHayL as well as the infrastructures themselves, in preparation for Sec. V, which compares data from the original IllinoisGRMHD code and various new implementations using GRHayL. Finally, in Section VI we summarize the library features and discuss planned extensions of GRHayL’s capabilities and implementations in additional infrastructures. In addition, we include two appendixes. Appendix A briefly describes the GRMHD evolution equations and our choice of variables used for evolving them. Appendix B details the automated, function-level continuous-integration testing that we perform to validate the library for each change made to the code repository.
@@ II Library Features
The core part of the library—the “chalice” of GRHayL—is the glue that ties all the separate modules together. This must be as minimal as possible to meet the aforementioned requirements of the library. GRHayL’s core contains definitions of C structs and simple helper functions for packing and unpacking data to and from these structs, as well as performing basic operations dependent purely on struct quantities [e.g., computing $g_{\mu\nu}$ and $g^{\mu\nu}$ from the Arnowitt-Deser-Misner (ADM) variables].
@@ II.1 EOS Gem
The EOS gem provides EoS functionality within GRHayL. It currently supports hybrid piecewise polytropic EoS and includes infrastructure for fully tabulated EoS. Tabulated EoS support is presently under development and is not yet intended for production simulations. Many other gems depend on functions provided by the EOS gem, but this dependency has been abstracted to allow users to implement their own EoS driver without changing the other gems. The EoS information is conveyed primarily via an EOS struct, in addition to several function pointers that are set during initialization. Users are free to swap the EOS functions with their own by setting these function pointers. Further, other EoS types or implementations can be easily added to the gem. For example, providing support for the Singularity EOS library would simply require defining the proper wrapper functions that convert the expected arguments for the GRHayL functions and making the appropriate call to Singularity EOS.
A hybrid EoS [65] assumes that the pressure and specific internal energy can be split into cold and thermal components, i.e.,
@@
In GRHayL, the cold pressure is represented by a piecewise polytrope (see, e.g., [96]),
@@
where $\Gamma_{\mathrm{th}}$ is a constant parameter that determines the conversion efficiency of kinetic to thermal energy at shocks. The EOS gem additionally supports a special case of the hybrid EoS called the simple EoS. [^2] This is often referred to as the ideal fluid, ideal gas, or Gamma-law EoS. In this case, the EoS contains a single piece with $\Gamma_{0}=\Gamma_{\mathrm{th}}$, resulting in $P=(\Gamma_{0}-1)\rho\epsilon$.
@@ II.2 Con2Prim Gem
The Con2Prim gem provides several methods for recovering the primitive variables $\bm{P}$ from the conservative variables $\bm{C}$ (defined in Eqs. 12 and 16, respectively). Because of the nonlinearity of the equations, a root-finding method such as Newton–Raphson is required. In addition, this gem provides several helper functions to compute $\bm{C}(\bm{P})$, to produce initial guesses for the primitives, and to enforce bounds on $\bm{C}$ [47, 43] and $\bm{P}$.
The gem includes several primitive recovery routines, including some by Noble et al. [86, 87], Font et al. [49, 42], Palenzuela et al. [91], and ongoing efforts to support Newman & Hamlin [85] (see also [115] for an entropy variation of the last two). Routines are named based on the first author of their originating paper and the dimensionality of the method (e.g., Noble2D, Palenzuela1D, etc.). The gem also provides several useful diagnostics, including the number of iterations a routine took to converge to the solution and whether or not backup routines were used.
The routines are adapted from origIGM and from [108], noting that origIGM originally adapted the Noble et al. routines from HARM3D [86, 87] and HARM3D+NUC [83]. We summarize all available primitive recovery routines currently implemented in GRHayL in Table 1.
@@
Noble2D [86]
Noble1D [86]
Noble1D_entropy [87]
Font1D [49, 42]
Palenzuela1D [91]
Palenzuela1D_entropy [115]
@@ II.3 Flux_Source Gem
The Flux_Source gem provides functions to compute the flux and source terms for the right-hand sides (RHSs) of the evolved GRMHD variables as described in Appendix A. These functions calculate the flux contributions to the RHSs using a finite-volume method, and the needed reconstructions can be performed using the Reconstruction gem. Flux_Source uses NRPy [100] to automatically generate optimized C code from the basic flux and source equations. The Flux_Source gem currently only provides the Harten-Lax-van Leer (HLL) [61] approximate Riemann solver, which has the form
@@
Here $c^{-}$ is the minimum characteristic speed, $c^{+}$ is the maximum characteristic speed, $F_{r,l}$ are the hydrodynamic fluxes defined in Eq. (20), and $U_{r,l}$ are the reconstructed values for the conserved variable $U$. The $r,l$ subscripts indicate fluid and spacetime quantities reconstructed at the right and left faces along some $i$-th flux direction. To compute the characteristic speeds, we follow the prescription given by [55], which approximates the general GRMHD dispersion relation as a quadratic in the wave speed. The wave speeds are then given by evaluating the quadratic equations